If I make $7.11 per hr and I work 4 hr a day and their are 168 hr a week how can i figer out what i will make
1
Expert's answer
2013-03-01T03:45:01-0500
We have a total number of hours per week - 168 hr.
Number of working hours - 4 hr a day.
The resulting earnings per hour of work - $7.11 per hr.
We can determine the earnings for the day, it equals $7.11×4=$28.44.
If we consider the five-working day week (number of working hours is equal to 20hr), then the earnings which will be received = $7.11×4×5=$142.2
If we consider the seven-working day week (number of working hours is equal to 28hr), then the earnings which will be received = $7.11×4×7=$199.08